For a journey, if the train increases its speed by 2km/h, then it will take 4 hours less to finish the trip. If it reduces its speed by 2km/h, then it will take 6 hours more. Find the distance that the train has to travel
Answer by mananth(16946)

let x be the distance traveled
y the speed
x/y = t the right time
Increased speed
x/(y+2) = t-4
decreased speed
x/(y-2)=t+6
x/y =t
x=yt
substitute x in (1)
yt/(y+2) = t-4
yt=(y+2)(t-4)
yt=yt-4y+2t-8
2t-4y=8..........................(1)
--
x/(y-2)=t+6
yt/(y-2)=t+6
yt=(y-2)(t+6)
yt=yt+6y-2t-12
2t-6y=-12...............................(2)
2 t -4 y = 8 .............1
Total value
2 t -6 y = -12 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
-6 t 12 y = -24
4 t -12 y = -24
Add the two equations
-2 t = -48
/ -2
t = 24
plug value of t in (1)
2 t -4 y = 8
48 -4 y = 8
-4 y = 8 -48
-4 y = -40
y = 10
x=yt
x=24*10
=240 miles